# Working Paper: Active Causal Structure Learning with Latent Variables: Towards Learning to Detour in Autonomous Robots

### 12 Mar 2026 v2

@@ 2.4.1 Problem definition and challenges
Two DAGs H, $K$ are $I$-equivalent if they have the same set of nodes $\mathcal{V}$ and represent the same conditional independence relations, i.e
@@
    $G$ is an $I$-map of $P$ if all conditional independence relations shown in $G$ occur in $P$, but all those in $P$ are not necessarily shown in $P$.
@@
    $G$ is a perfect map of $P$ if $G$ is a $D$-map and $I$-map of $P$ at the same time, meaning that all relations of $P$ are shown in $G$, and all relations of $G$ are given in $P$.
@@ 2.4.2 Structural learning with latent variables
  • A latent variable $LV$, is a variable unobserved by the agent and placed naturally in the environment, whether or not it is relevant to the problem faced by the agent.
  • A hidden variable $HV$, is considered to be a representation of one or several latent variables generated by the agent with the intention of modeling the latent variables. Since the latent variables cannot be fully observed, the agent needs to model or estimate them. Generally, the agent will only want to model those latent variables that are relevant to find a better solution to the problem the agent faces.
@@ 3.3 Actions and Observations Space
  • BarrierTactileBarrier\;Tactile: This is a binary variable that is $1$ if the infinity norm between the center of the agent and any of the spikes of the barrier is lower than $2$, and 0 otherwise.
@@
  • TargetinvisualFieldTarget\;in\;visual\;Field: A binary variable, it will be $1$ if $Depth$ is lower than $2$ and HeadingangleHeading\;angle is between $\frac{\pi}{2},\frac{\pi}{2}]$.
@@ 3.4 Agent´s initial Dynamic Decision Network
where $sf$ and $sa$ refer to the decision values that the actions StepForwardStep\;Forward and StepAsideStep\;Aside take, respectively.
@@ 3.5 Problem modeled as a POMDP
  • $T$ would be the degenerate distribution, which gives probability $1$ to the new state $s^{\prime}$ determined by the mechanistic function of the executed action $a$, according to the state $s$ of the agent.
@@
  • $O$ would also be the degenerate distribution, assigning probability $1$ to the value given by the function of the observation variable. In other words, the agent has accurate sensors that do not introduce errors into the measurement process.
@@ 4.1 Causal Discovery
This normalization factor results from working with percentages of entropy reduction relative to the entropy of the $Obs$ variable alone. The greater the reduction in entropy, the more information the variables contribute to the explanation of the $Obs$ variable. In this regard, the algorithm will work as a forward algorithm, that is the observation variables are added, apart from the action variable, according to the value of the partial entropy reduction; if the partial reduction is not enough, the process of adding variables is stopped.
@@ 5.1.1 Utility Surprise
Case $1$ means that the utility value obtained had a very low associated probability, and if, in addition, the utility value is lower than the mean, then it means that there was some unobserved variable that influenced the agent’s action, causing the utility not to be maximized as expected, and making it convenient for the agent to learn to predict when something similar might occur so that the decision is correct. The reason for studying whether the expected value is lower or higher is that, in the end, the agent’s choice of decisions is conditioned to look for the highest expected value. So, if it turns out that the value obtained is lower, then the agent could have chosen another decision that might have yielded a higher utility. It is also irrelevant whether the dispersion of the distribution was high or low, since the surprise coefficient indicates whether the utility obtained was a surprise event or not.
As for the case $2$, if the utility obtained is higher than expected, then we would have the opposite case, i.e. it has been discovered that there is a variable that influences and helps to obtain higher utility values, so it is convenient to learn to model when this may occur. However, in this paper, we will focus on detecting the latent variables that have a negative influence on the utility, and therefore the probability of influence will be close to 0, in the case that the utility obtained is higher than expected.
If the surprise coefficient is close to 0, as for the cases $3$ and $4$, then it means that the probability that there is a variable that the agent does not observe, and that changes its interaction with the environment will be low, since the agent is able to correctly predict the values of the observation variables that it will get after executing the actions. Taking all this into account, the coefficient that will be responsible for making this comparison will be an extrapolation of the one described in (22), which will also include whether the utility obtained is higher or lower than the one expected.
@@ 5.1.2 Probability of Influence of the Hidden Variable
The probability of influence of the latent variable will be a function dependent on the value of the coefficient of the utility surprise. The range of possible values that $C_{U}$ can take is potentially $(-\infty,\infty)$, and the goal is to convert it to the range $[0,1]$. Therefore, assuming we want to calculate $\text{I\kern-1.49994ptP}(HV=0)$, the lower the coefficient, the greater the negative surprise, that is, the greater the evidence that $HV=1$, so $\text{I\kern-1.49994ptP}(HV=0)$ must tend to 0 . On the other hand, the higher it is, the greater the evidence that $HV=0$, thus $\text{I\kern-1.49994ptP}(HV=0)$ must tend to $1$. However, when $C_{U}$ is close to 0, this will indicate that there has not been enough surprise at instant t, so we do not have enough evidence to say that $HV=1$, but neither is it correct to say that $HV=0$, since it could happen that the influence of the latent variable is practically null, therefore, when $C_{U}=0$ we will have that $\text{I\kern-1.49994ptP}(HV=0)=\frac{1}{2}$.
@@
which behaves as we have previously described. That is, when $C_{U}<<0$, the probability of no influence will tend to 0; meanwhile if $C_{U}>>0$, the probability will tend to $1$.
@@ 5.1.3 Selection of related variables: Surprise in Observations
Returning to the Learning to Detour case (Figure 5), when the agent first bumps into the barrier, an unexpected effect on the BarrierTactileBarrier\;Tactile Observation variable occurs, since it is not predicted by any existing transition. Similarly, an unexpected effect on the $Depth$ Observation variable occurs after the activation of the StepForwardStep\;Forward Action variable, since the associated transition predicts a reduction on the $Depth$ value, yet the $Depth$ value remains unaltered since no forward movement can take place, due to the blocking effect by the barrier. For the purposes of visualization, the surprises were normalized by dividing them by the sum of the surprises present in the observed variables. This allowed for the accurate representation (Figure 5) of the fact that the variables BarrierTactileBarrier\;Tactile and $Depth$ experienced the greatest degree of surprise, and therefore represent the variables that should be selected.
@@ 5.1.4 Structure Learning with Hidden Variable
    As we previously mentioned, the hidden variable at time $t$ can not be observed directly, yet we can assign a probability distribution based on the values of the $Obs_{t}$ related with the latent variable. This means that, based on the mechanistic causality definition, there exists a function such that the $HV$ values can be generated from the values of $Obs_{t}$ and a random noise, so there must be edges from $Obs_{t}$ to $HV$.
@@
    On the other hand, as discussed in section 5.1.2, the latent variable influences the probability distribution of the observed variables at $t+1$, causing a surprise observation at $Obs_{t+1}$. Therefore, the latent variable influences the generation of the $Obs_{t+1}$ values, and consequently, the latent variable is part of the input mechanistic function. Since the hidden variable must model the causal relations of the latent variable, there must be edges from $HV$ to $Obs_{t+1}$.
@@
    Additionally, the edges from $Obs_{t}$ to $Obs_{t+1}$ are kept since the presence of $HV$ does not replace the causal relations previously learned by the agent, but conditions them.
@@ 5.2 CPT Estimation
Returning to the Learning to Detour case, after learning, the hidden variable $HV$, a new transition must be learned so that expected observation $BT=1$ is realized when the Step Forward action is activated very close to the barrier. Also, after learning the hidden variable $HV$, a new transition must be learned so that the expected observation in $Depth$ is realized when the Step Forward action is activated and the agent is close to the barrier. Lastly, when the agent is close to the end of the barrier, the expected observation of $Tactile=0$ is realized, when Step Aside is selected to the right.
@@ 5.3 Learning to Detour behavior
Thus, after learning the hidden variable $HV$, it can be observed from Figure 8. that the agent’s behaviour has undergone an important transformation compared to the behaviour before learning.
@@
The first thing to notice, is the shift from the agent’s exclusive reliance on the StepForwardStep\;Forward action (before learning) to a more nuanced approach (during learning), involving a reduction in the power of StepForwardStep\;Forward decisions and the initiation of StepAsideStep\;Aside actions to the right. This is due to the changes made to the CPT of $BT_{t+1}$. As can be seen in the marginalized CPT of $BT_{t+1}$ before and after learning (Table 2), there is a greater probability of a $BT=1$ in the presence of the $HV$. This CPT, along with the CPT learned from $HV$ (Table 3), enables the agent to act in this way.
@@
The second significant event occurs when the agent collides with the barrier. During learning, the agent reduces the power of the StepForwardStep\;Forward action. This is a consequence of the agent’s adaptation, which has led to a reduction in the expected utility of activating StepForwardStep\;Forward when the probability of $HV=1$ is high. It is important to notice that at this moment, another important change takes place within our framework. That is, one of the key objectives is to enable the agent to learn from unexpected outcomes, which will enable it to make more accurate predictions and, therefore, avoid surprises in the future. In our case, Figures 9 and 10 illustrate that the surprises associated with $BT$ (Barrier Tactile) and $D$ (Depth) are reduced in comparison to those observed prior to learning. This clearly indicates that the predictions made by the agent are more accurate with respect to the environment when the $HV$ factor is present. In summary, before learning, the agent tends to go straight to the barrier and bumps into it in many occasions. Yet, after learning, the agent tends to avoid the barrier by side stepping around it in its trajectory to the target.
@@
Finally, after overcoming the obstacle, the agent makes essentially the same actions in both scenarios, demonstrating that the incorporation of the $HV$ and the acquisition of CPTs did not impede the agent’s appropriate behaviour when the probability of $HV=1$ is low, allowing the agent to successfully reach the target.
@@ 0.B.1.1 Actions Space
To go from discrete to continuous values it has been decided to select a random decision value using a uniform distribution with the appropriate bounds, depending on the discrete decision selected by the agent. For instance, if the agent decide to do StepForwardStep\;Forward with a value $1$, then it will returned to the Environment to execute $SA(x,y,\alpha,s)$ with $s\sim U(0.5,1)$.
@@ 0.B.1.2 Observation Space
In the Observation Space, $Depth$ and HeadingAngleHeading\;Angle are the only observation variables that need a discretization.
For $Depth$, the interval $0,d_{max}]$ has been divided into five equal parts, where $d_{max}=\sqrt{x_{max}^{2}+y_{max}^{2}}$
@@ 0.B.2 How the Agent and Environment interacts: Restrictors
The environment will have two restrictors: $MapBounds$ and $BarrierImpact$.
@@ 0.B.2.1 MapBounds
The first restrictor ensures that the agent cannot exceed the limits set by the environment, that said, the agent is unable to perform an action that would result in its displacement outside of the rectangle $ABCD$, with $A=(0,0)$, $B=(10,0)$, $C=(10,15)$, $D=(0,15)$.
@@ 0.B.2.2 BarrierImpact
On the other hand, the $BarrierImpact$ restrictor works in a similar way to the previous one, but with some nuances. This restrictor will detect when an action performed by the agent causes it to go beyond the barrier, impacting with any of its spikes. In this instance, rather than denying the action and preventing its execution, the value of the action will be substituted with the maximum value that does not activate the restrictor.
@@ 0.B.3 Transition Probabilities
This section outlines the transition probability matrices employed by the agent. Each matrix consists of rows representing the values of $Obs_{t}$ and columns representing the values of $Obs_{t+1}$. As each probability transition $Obs_{t+1}$ may depends on the value of StepForwardStep\;Forward and StepAsideStep\;Aside, the matrices are expressed in terms of the value of $SF$ and $SA$.
@@ 0.B.3.2 HeadingAngleHeading\;Angle
The parameter $p$ serves to balance the importance of $SA$ and $SF$ in the transition probability. After testing that the agent behaved correctly in different scenarios the parameter was set to $p=0.2$

### 26 Mar 2026 v3

@@ 2.4.1 Problem definition and challenges
Two DAGs H, $K$ are II-equivalent if they have the same set of nodes $\mathcal{V}$ and represent the same conditional independence relations, i.e
@@
    $G$ is an II-map of $P$ if all conditional independence relations shown in $G$ occur in $P$, but all those in $P$ are not necessarily shown in $P$.
@@
    $G$ is a perfect map of $P$ if $G$ is a $D$-map and II-map of $P$ at the same time, meaning that all relations of $P$ are shown in $G$, and all relations of $G$ are given in $P$.
@@ 2.4.2 Structural learning with latent variables
  • A latent variable LVLV, is a variable unobserved by the agent and placed naturally in the environment, whether or not it is relevant to the problem faced by the agent.
  • A hidden variable HVHV, is considered to be a representation of one or several latent variables generated by the agent with the intention of modeling the latent variables. Since the latent variables cannot be fully observed, the agent needs to model or estimate them. Generally, the agent will only want to model those latent variables that are relevant to find a better solution to the problem the agent faces.
@@ 3.3 Actions and Observations Space
  • BarrierTactileBarrier\;Tactile: This is a binary variable that is 11 if the infinity norm between the center of the agent and any of the spikes of the barrier is lower than 22, and 0 otherwise.
@@
  • TargetinvisualFieldTarget\;in\;visual\;Field: A binary variable, it will be 11 if DepthDepth is lower than 22 and HeadingangleHeading\;angle is between $\frac{\pi}{2},\frac{\pi}{2}]$.
@@ 3.4 Agent´s initial Dynamic Decision Network
where sfsf and sasa refer to the decision values that the actions StepForwardStep\;Forward and StepAsideStep\;Aside take, respectively.
@@ 3.5 Problem modeled as a POMDP
  • $T$ would be the degenerate distribution, which gives probability 11 to the new state $s^{\prime}$ determined by the mechanistic function of the executed action $a$, according to the state $s$ of the agent.
@@
  • $O$ would also be the degenerate distribution, assigning probability 11 to the value given by the function of the observation variable. In other words, the agent has accurate sensors that do not introduce errors into the measurement process.
@@ 4.1 Causal Discovery
This normalization factor results from working with percentages of entropy reduction relative to the entropy of the ObsObs variable alone. The greater the reduction in entropy, the more information the variables contribute to the explanation of the ObsObs variable. In this regard, the algorithm will work as a forward algorithm, that is the observation variables are added, apart from the action variable, according to the value of the partial entropy reduction; if the partial reduction is not enough, the process of adding variables is stopped.
@@ 5.1.1 Utility Surprise
Case 11 means that the utility value obtained had a very low associated probability, and if, in addition, the utility value is lower than the mean, then it means that there was some unobserved variable that influenced the agent’s action, causing the utility not to be maximized as expected, and making it convenient for the agent to learn to predict when something similar might occur so that the decision is correct. The reason for studying whether the expected value is lower or higher is that, in the end, the agent’s choice of decisions is conditioned to look for the highest expected value. So, if it turns out that the value obtained is lower, then the agent could have chosen another decision that might have yielded a higher utility. It is also irrelevant whether the dispersion of the distribution was high or low, since the surprise coefficient indicates whether the utility obtained was a surprise event or not.
As for the case 22, if the utility obtained is higher than expected, then we would have the opposite case, i.e. it has been discovered that there is a variable that influences and helps to obtain higher utility values, so it is convenient to learn to model when this may occur. However, in this paper, we will focus on detecting the latent variables that have a negative influence on the utility, and therefore the probability of influence will be close to 0, in the case that the utility obtained is higher than expected.
If the surprise coefficient is close to 0, as for the cases 33 and 44, then it means that the probability that there is a variable that the agent does not observe, and that changes its interaction with the environment will be low, since the agent is able to correctly predict the values of the observation variables that it will get after executing the actions. Taking all this into account, the coefficient that will be responsible for making this comparison will be an extrapolation of the one described in (22), which will also include whether the utility obtained is higher or lower than the one expected.
@@ 5.1.2 Probability of Influence of the Hidden Variable
The probability of influence of the latent variable will be a function dependent on the value of the coefficient of the utility surprise. The range of possible values that $C_{U}$ can take is potentially $(-\infty,\infty)$, and the goal is to convert it to the range $[0,1]$. Therefore, assuming we want to calculate $\text{I\kern-1.49994ptP}(HV=0)$, the lower the coefficient, the greater the negative surprise, that is, the greater the evidence that $HV=1$, so $\text{I\kern-1.49994ptP}(HV=0)$ must tend to 0 . On the other hand, the higher it is, the greater the evidence that $HV=0$, thus $\text{I\kern-1.49994ptP}(HV=0)$ must tend to 11. However, when $C_{U}$ is close to 0, this will indicate that there has not been enough surprise at instant t, so we do not have enough evidence to say that $HV=1$, but neither is it correct to say that $HV=0$, since it could happen that the influence of the latent variable is practically null, therefore, when $C_{U}=0$ we will have that $\text{I\kern-1.49994ptP}(HV=0)=\frac{1}{2}$.
@@
which behaves as we have previously described. That is, when $C_{U}<<0$, the probability of no influence will tend to 0; meanwhile if $C_{U}>>0$, the probability will tend to 11.
@@ 5.1.3 Selection of related variables: Surprise in Observations
Returning to the Learning to Detour case (Figure 5), when the agent first bumps into the barrier, an unexpected effect on the BarrierTactileBarrier\;Tactile Observation variable occurs, since it is not predicted by any existing transition. Similarly, an unexpected effect on the DepthDepth Observation variable occurs after the activation of the StepForwardStep\;Forward Action variable, since the associated transition predicts a reduction on the DepthDepth value, yet the DepthDepth value remains unaltered since no forward movement can take place, due to the blocking effect by the barrier. For the purposes of visualization, the surprises were normalized by dividing them by the sum of the surprises present in the observed variables. This allowed for the accurate representation (Figure 5) of the fact that the variables BarrierTactileBarrier\;Tactile and DepthDepth experienced the greatest degree of surprise, and therefore represent the variables that should be selected.
@@ 5.1.4 Structure Learning with Hidden Variable
    As we previously mentioned, the hidden variable at time $t$ can not be observed directly, yet we can assign a probability distribution based on the values of the $Obs_{t}$ related with the latent variable. This means that, based on the mechanistic causality definition, there exists a function such that the HVHV values can be generated from the values of $Obs_{t}$ and a random noise, so there must be edges from $Obs_{t}$ to HVHV.
@@
    On the other hand, as discussed in section 5.1.2, the latent variable influences the probability distribution of the observed variables at $t+1$, causing a surprise observation at $Obs_{t+1}$. Therefore, the latent variable influences the generation of the $Obs_{t+1}$ values, and consequently, the latent variable is part of the input mechanistic function. Since the hidden variable must model the causal relations of the latent variable, there must be edges from HVHV to $Obs_{t+1}$.
@@
    Additionally, the edges from $Obs_{t}$ to $Obs_{t+1}$ are kept since the presence of HVHV does not replace the causal relations previously learned by the agent, but conditions them.
@@ 5.2 CPT Estimation
Returning to the Learning to Detour case, after learning, the hidden variable HVHV, a new transition must be learned so that expected observation $BT=1$ is realized when the Step Forward action is activated very close to the barrier. Also, after learning the hidden variable HVHV, a new transition must be learned so that the expected observation in DepthDepth is realized when the Step Forward action is activated and the agent is close to the barrier. Lastly, when the agent is close to the end of the barrier, the expected observation of $Tactile=0$ is realized, when Step Aside is selected to the right.
@@ 5.3 Learning to Detour behavior
Thus, after learning the hidden variable HVHV, it can be observed from Figure 8. that the agent’s behaviour has undergone an important transformation compared to the behaviour before learning.
@@
The first thing to notice, is the shift from the agent’s exclusive reliance on the StepForwardStep\;Forward action (before learning) to a more nuanced approach (during learning), involving a reduction in the power of StepForwardStep\;Forward decisions and the initiation of StepAsideStep\;Aside actions to the right. This is due to the changes made to the CPT of $BT_{t+1}$. As can be seen in the marginalized CPT of $BT_{t+1}$ before and after learning (Table 2), there is a greater probability of a $BT=1$ in the presence of the HVHV. This CPT, along with the CPT learned from HVHV (Table 3), enables the agent to act in this way.
@@
The second significant event occurs when the agent collides with the barrier. During learning, the agent reduces the power of the StepForwardStep\;Forward action. This is a consequence of the agent’s adaptation, which has led to a reduction in the expected utility of activating StepForwardStep\;Forward when the probability of $HV=1$ is high. It is important to notice that at this moment, another important change takes place within our framework. That is, one of the key objectives is to enable the agent to learn from unexpected outcomes, which will enable it to make more accurate predictions and, therefore, avoid surprises in the future. In our case, Figures 9 and 10 illustrate that the surprises associated with BTBT (Barrier Tactile) and $D$ (Depth) are reduced in comparison to those observed prior to learning. This clearly indicates that the predictions made by the agent are more accurate with respect to the environment when the HVHV factor is present. In summary, before learning, the agent tends to go straight to the barrier and bumps into it in many occasions. Yet, after learning, the agent tends to avoid the barrier by side stepping around it in its trajectory to the target.
@@
Finally, after overcoming the obstacle, the agent makes essentially the same actions in both scenarios, demonstrating that the incorporation of the HVHV and the acquisition of CPTs did not impede the agent’s appropriate behaviour when the probability of $HV=1$ is low, allowing the agent to successfully reach the target.
@@ 0.B.1.1 Actions Space
To go from discrete to continuous values it has been decided to select a random decision value using a uniform distribution with the appropriate bounds, depending on the discrete decision selected by the agent. For instance, if the agent decide to do StepForwardStep\;Forward with a value 11, then it will returned to the Environment to execute $SA(x,y,\alpha,s)$ with $s\sim U(0.5,1)$.
@@ 0.B.1.2 Observation Space
In the Observation Space, DepthDepth and HeadingAngleHeading\;Angle are the only observation variables that need a discretization.
For DepthDepth, the interval $0,d_{max}]$ has been divided into five equal parts, where $d_{max}=\sqrt{x_{max}^{2}+y_{max}^{2}}$
@@ 0.B.2 How the Agent and Environment interacts: Restrictors
The environment will have two restrictors: MapBoundsMapBounds and BarrierImpactBarrierImpact.
@@ 0.B.2.1 MapBounds
The first restrictor ensures that the agent cannot exceed the limits set by the environment, that said, the agent is unable to perform an action that would result in its displacement outside of the rectangle ABCDABCD, with $A=(0,0)$, $B=(10,0)$, $C=(10,15)$, $D=(0,15)$.
@@ 0.B.2.2 BarrierImpact
On the other hand, the BarrierImpactBarrierImpact restrictor works in a similar way to the previous one, but with some nuances. This restrictor will detect when an action performed by the agent causes it to go beyond the barrier, impacting with any of its spikes. In this instance, rather than denying the action and preventing its execution, the value of the action will be substituted with the maximum value that does not activate the restrictor.
@@ 0.B.3 Transition Probabilities
This section outlines the transition probability matrices employed by the agent. Each matrix consists of rows representing the values of $Obs_{t}$ and columns representing the values of $Obs_{t+1}$. As each probability transition $Obs_{t+1}$ may depends on the value of StepForwardStep\;Forward and StepAsideStep\;Aside, the matrices are expressed in terms of the value of SFSF and SASA.
@@ 0.B.3.2 HeadingAngleHeading\;Angle
The parameter $p$ serves to balance the importance of SASA and SFSF in the transition probability. After testing that the agent behaved correctly in different scenarios the parameter was set to $p=0.2$